if you halve my age and add on 21, you will get three-quarters of the age that i was 4 years ago.
Answer by Boreal(15235)

age now is x
(x/2)+21=(3/4)(x-4)
multiply everything by 4 to clear fractions
2x+84=3(x-4)=3x-12
x=96 by moving terms
96 years old ANSWER
half is 48 and add 21 is 69.
4 years ago, you were 92, and 3/4 of that is 69.